If the inverse demand curve of profit maximizing monopolist is given as P =1200 − 2Q , and cost function as 

 C = Q3 − 61.25Q2+1528.5Q + 2000, find equilibrium output level, monopolist price, and profit.


1
Expert's answer
2021-07-27T16:36:02-0400

"P= 1200-2Q"

"C= Q^3-61.25Q^2+1528.5Q" +2000

Total Revenue= Price ×Quantity ("P\u00d7Q)"

Therefore TR = "(1200-2Q)Q" = 1200Q-2Q2


Marginal Revenue"= \frac{\delta TR}{\delta Q}" =

"1200- 4Q"


Marginal Cost "= \frac{\delta TC}{\delta Q}" =

"3Q^2 - 122.5Q+1528.5"


At equilibrium output "MR=MC"

"1200- 4Q =3Q^2-122.5Q+1528.5\implies"

"3Q^2-118.5Q+328.5=0"

"Q= 36.5" or 3

Equilibrium output= "36.5"


Using Q = 36.5, the monopolist price is

"P= 1200-2(36.5) = 1127"


Profit is obtained by getting the difference between total revenue and the total cost.

Profit = "TR- TC"

"TR = 1200(36.5)-2(36.5)^2"

"= 41135.5"

".TC = 36.5^3- 61.25(36.5^2)" + 1528.5(36.5)+2000 = 24817.0625


Profit = "41135.5- 24817.0625" = 16318.4375